# Activation Function Design Sustains Plasticity in Continual Learning

### v1

@@ 1 Introduction
  • Two drop-in activation functionsSmooth-Leaky and Randomized Smooth-Leaky—that preserve a non-zero derivative floor and target the moderate-leak regime with a $C^{1}$ transition, improving continual adaptation without adding parameters (Secs. 6, 7).
@@ 2 Activation Functions and Plasticity in Continual Learning
Rectifiers. Classic ReLU is efficient but prone to the dead‑unit problem (Maas et al., 2013): neurons that output 0 receive no gradient and thus become inactive permanently. Continual‑learning studies confirm a rising fraction of dormant units over time (dormant neurons phenomenon Sokar et al. (2023)), shrinking gradient norms and eventually freezing learning (Abbas et al., 2023; Dohare et al., 2024). ReLU networks often show an increase in parameter norms during training (as they drive outputs with ever-larger weights) and a drop in the number of effective directions in weight space that can reduce error Lewandowski et al. (2023). Leaky‑ReLU alleviates this with a constant negative slope, while PReLU (He et al., 2015) and RReLU (Xu et al., 2015) postpone dormancy by making the negative slope learnable or by randomly sampling it during training.
Saturating sigmoids. Sigmoid and Tanh map inputs to bounded ranges; when units saturate, derivatives shrink toward zero and gradients can vanish, slowing learning Glorot & Bengio (2010). This is relevant for continual learning, which requires sustained adaptation under shift and where loss of plasticity has been repeatedly observed Dohare et al. (2024); Abbas et al. (2023); Juliani & Ash (2024). (See Sec. 4 for empirical evidence.)
Smooth non-monotonic. Swish (Ramachandran et al., 2017) and GeLU (Hendrycks & Gimpel, 2016) are smooth, weakly non-monotonic rectifiers that preserve small—but non-zero—gradients for inputs near and below zero. This mitigates “dying ReLU” behavior (Maas et al., 2013), so units remain trainable when pre-activation distributions drift under shift. In continual settings, this responsiveness supports ongoing adaptation; empirically, non-monotonic/smooth rectifiers have shown advantages in both supervised and RL domains (Ramachandran et al., 2017; Hendrycks & Gimpel, 2016; Elfwing et al., 2018), and in our experiments (Secs. 34) they exhibit lower dead-unit fractions and stronger late-cycle adaptation than zero-floor rectifiers.
@@
Under identical models and training budgets on Split-CIFAR-100, activation function rankings compress in i.i.d. joint training but separate sharply in class-incremental (C-IL) settings (Van de Ven et al., 2022) (see Table 1). This motivates probing how negative-branch behavior affects plasticity under shift. Unless noted, all case studies use the same 4-layer CNN backbone, optimizer, and training budget (full details in App. B), isolating activation effects from architectural or optimization confusion.
@@
i.i.d. 72.11 72.00 73.71 71.43 73.16 72.51 72.66 72.64 72.12 66.49 58.78
C-IL 24.41 28.57 32.95 22.71 24.43 20.91 22.79 27.59 27.49 26.44 25.47
Table 1: Split-CIFAR-100 average accuracy (5 runs) with identical architecture, optimizer, and budget. Performance differences across activation functions are modest under i.i.d. joint training but widen under class-incremental learning (C-IL). RReLU attains the top mean in both settings; the C-IL improvement is statistically significant (all $p<0.05$), whereas i.i.d. differences are not.
@@ 3 Case Study 1: Negative-Slope ‘Goldilocks Zone’
‘Goldilocks zone’ for negative slopes (with shape-matched comparison). For activations with a constant negative-branch slope (Leaky-ReLU, RReLU), performance reliably peaks for a moderate leak $0.6\!\lesssim\!\bar{s}\!\lesssim\!0.9$ and degrades once $\bar{s}\!\gtrsim\!0.9$ (Fig. C1, Tab. 1), suggesting the existence of a ‘Goldilocks zone’ for negative slopes. To compare smooth-tailed functions (ELU/CELU/SELU, Swish, GELU) on a common scale, we project their negative-branch behavior onto an effective slope axis, $\bar{s}=\mathbb{E}_{x<0}[\varphi^{\prime}(x)]$, representing the average derivative for negative inputs. When matched by $\bar{s}$, these smooth tails still underperform linear leaks within the ‘Goldilocks zone’ and exhibit higher dead-unit fractions, approaching (but not surpassing) the linear-leak peak only for $\bar{s}>1$ (App. C.1).
Failure Modes for Slope Magnitude. Building on the ‘Goldilocks zone’ phenomenon, we examined why performance degrades outside it. Across activations we consistently observe a dead-unit regime as $\bar{s}\!\to\!0$ ($\approx\!45\%$), and accuracy tracks this effect closely: Final Average Accuracy (ACC) (Sec. B.5.1) is moderately and significantly correlated with dead units (Pearson (r)=0.59(r){=}\,-0.59, $p{=}5.7{\times}10^{-8}$), whereas its correlation with final scaled gradient norms is weaker but also significant (Pearson (r)=0.28(r){=}\,-0.28, p=0.019p{=}0.019) (App. C.2). At very large $\bar{s}$ (near/above $1$), we sometimes see conditioning spikes—higher dominant Hessian eigenvalue and lower gradient-Gram effective rank ($\lambda_{\max}\!\uparrow$, Effective rank $\downarrow$)— but these appear mainly for smooth-tailed shapes and are not universal for constant-slope leaks (Fig. C2; App. B.6, C.2). Therefore, dead units are the primary driver of failure, and we treat curvature as a useful diagnostic rather than a general mechanism, and focus our designs on maintaining negative-side responsiveness.
Current adaptive, learnable slopes fail and need constraints to stay in-band. Under non-stationarity, the “right” negative-side responsiveness is not uniform across units or tasks. We therefore asked whether learnable or randomized slopes can find and maintain their specific ‘Goldilocks zone’. Per-neuron PReLU (PReLU-N) drifts below the band ($\approx 0.3-0.6$) over training and attains 30.1% ACC (see Figs.. C.3); layer/global scopes drift even further (Figs. C5). Thus, adaptivity is relevant—it offers robustness when a single fixed leak cannot serve all units—but unconstrained adaptation does not reliably remain in the pre-defined ‘Goldilocks zone’, which explains why learned slopes may sit outside it despite good (but suboptimal) ACC.
@@ 4 Case Study 2: Desaturation Dynamics Under Shocks
Every $C_{\ell}{=}10$ epochs we apply a one-epoch scaling shock by multiplying all pre-activations $z$ by $\gamma\in\{1.5,\,0.5,\,0.25,\,2.0\}$, then revert to $\gamma{=}1$ (App. D.1). Large $\gamma$ (e.g., $2.0$) pushes $z$ into negative tails or positive plateaus (saturation); small $\gamma$ (e.g., $0.5$) produces the mirrored event. Each activation uses its best negative-slope setting from Case Study 3 (Table B2). A unit is saturated at a step if $|\varphi^{\prime}(x)|<10^{-3}$. We report: (i) Peak SF: the maximum saturated fraction immediately after each shock; (ii) AUSC: the area under the saturation curve over the recovery window (lower is better); (iii) Recovery time $\bm{\tau_{95}}$: steps needed to regain $95\%$ of pre-shock performance (App. D.5).
@@
The width of the dead band predicts shock sensitivity. Beyond the findings of derivative floor and sidedness, we ask how much of the input range of an activation produces nearly zero gradients. We define a Dead-Band Width (DBW) as the fraction of a typical pre-activation range (e.g., [-100, 100]) where the magnitude of the activation’s first derivative, $\lvert\varphi^{\prime}(x)\rvert$ falls below certain threshold, $\epsilon<10^{-3}$. Analytically computed DBW, $|\varphi^{\prime}(x)|<10^{-3}$, strongly tracks desaturation outcomes across activations: it correlates with AUSC (r=0.81,p=0.0016r=0.81,\;p=0.0016) and with non-recovery (r=0.84,p=0.0013r=0.84,\;p=0.0013). Intuitively, wider dead bands expose more inputs to vanishing gradients during shocks, prolonging saturation and increasing failure rates (see App. D.4).
@@ 5.1 Smooth-Leaky and Randomized Smooth-Leaky
Figure 3: Smooth-Leaky with $\alpha{=}0.1$, p=3.0p{=}3.0, c=5.0c{=}5.0. Randomized Smooth-Leaky draws $\alpha$ from bounds; visually it matches Smooth-Leaky for the sampled $\alpha$.
@@
This randomized variant preserves the strict floor and $C^{1}$ transition while encouraging robustness to small variations in negative-side responsiveness. Limitations of multi-parameter design are explained in App. B.1.
@@ 6 Continual Supervised Learning
ReLU 78.85 20.03 25.79 4.76 73.71
Leaky-ReLU 84.14 91.53 98.34 48.86 85.28
Sigmoid 76.96 79.59 52.24 1.79 63.89
Tanh 70.32 63.40 58.56 28.59 70.97
RReLU 83.95 93.10 98.02 53.60 85.07
PReLU 82.62 92.67 96.86 43.30 82.37
Swish (SiLU) 83.41 67.73 87.40 35.31 82.64
GeLU 78.97 38.79 42.85 17.60 75.49
CeLU 82.93 37.16 29.64 54.23 81.14
eLU 80.50 84.23 57.45 47.64 80.10
SeLU 80.43 79.95 84.61 49.07 80.98
CReLU 82.66 89.47 92.90 20.56 84.85
Rational 80.18 91.27 94.82 40.41 80.65
Smooth-Leaky 84.03 91.69 98.36 49.87 85.38
Rand. Smooth-Leaky 84.26 93.33 98.42 57.01 86.23
Table 2: Total Average Online Task Accuracy (%) across all Continual Supervised Benchmark Problem by activation function. Averaged over 5 independent runs. Rand. Smooth-Leaky is statistically significant (p << 0.05) with respect to the next best-performing activation, Smooth-Leaky is also significant compared to the next in Rand. Label CIFAR, CIFAR 5+1 and Continual ImageNet.
Across the five continual benchmarks, we observe a clear pattern, reported in Tab. 2. First, rectifiers with a learnable or randomized negative branch dominate: Leaky-ReLU, RReLU, PReLU, Smooth-Leaky, and Rand. Smooth-Leaky consistently outperform ReLU—especially on the harder settings—while smooth rectifiers such as Swish/SiLU are competitive but typically trail the best leaky-family members. Second, we again observe the first reported ‘Goldilocks zone’ for the negative branch (cf. Sec. 3): the strongest performers cluster around an initial/effective negative slope in the range [0.6, 0.9][0.6,\,0.9] (including the mean of RReLU bounds and neuron-wise $\alpha$ in PReLU). We also evaluated CReLU and Rational activations, recently proposed to mitigate loss of plasticity Abbas et al. (2023); Kumar et al. (2023); Delfosse et al. (2021b). Constrained rationals were excluded, as previous work shows that they improve RL stability but reduce plasticity Surdej et al. (2025), which is our main focus. In our experiments, both CReLU and Rational outperform ReLU, but remain below other standard activations and our proposed variants (see Tab. E2 and Fig. E1).

### v2

@@ 1 Introduction
  • Two drop-in activation functionsSmooth-Leaky and Randomized Smooth-Leaky—that preserve a non-zero derivative floor and target the moderate-leak regime with a $C^{1}$ transition (i.e., having a continuous first derivative) between its linear and non-linear regions efficiently improving continual adaptation. (Secs. 6, 7).
@@ 2 Activation Functions and Plasticity in Continual Learning
Rectifiers. ReLU is efficient but prone to the dead‑unit problem (Maas et al., 2013): neurons that output 0 receive no gradient and thus become inactive permanently. Continual‑learning studies confirm a rising fraction of dormant units over time (dormant neurons phenomenon Sokar et al. (2023)), shrinking gradient norms and eventually freezing learning (Abbas et al., 2023; Dohare et al., 2024). ReLU networks often show an increase in parameter norms during training (as they drive outputs with ever-larger weights) and a drop in the number of effective directions in weight space that can reduce error Lewandowski et al. (2023). Leaky‑ReLU alleviates this with a constant negative slope, while PReLU (He et al., 2015) and RReLU (Xu et al., 2015) postpone dormancy by making the negative slope learnable or by randomly sampling it during training.
Saturating sigmoids. Sigmoid and Tanh map inputs to bounded ranges; when units saturate, derivatives shrink toward zero and gradients can vanish, slowing learning Glorot and Bengio (2010). This is relevant for continual learning, which requires sustained adaptation under shift and where loss of plasticity has been repeatedly observed Dohare et al. (2024); Abbas et al. (2023); Juliani and Ash (2024). (See Sec. 4 for empirical evidence.)
Smooth non-monotonic. Swish (Ramachandran et al., 2017) and GeLU (Hendrycks and Gimpel, 2016) are smooth, weakly non-monotonic rectifiers that preserve small—but non-zero—gradients for inputs near and below zero. This mitigates “dying ReLU” behavior (Maas et al., 2013), so units remain trainable when pre-activation distributions drift under shift. In continual settings, this responsiveness supports ongoing adaptation; empirically, non-monotonic/smooth rectifiers have shown advantages in both supervised and RL domains (Ramachandran et al., 2017; Hendrycks and Gimpel, 2016; Elfwing et al., 2018), and in our experiments (Secs. 34) they exhibit lower dead-unit fractions and stronger late-cycle adaptation than zero-floor rectifiers.
@@
Under identical models and training budgets on Split-CIFAR-100, activation function rankings compress in i.i.d. joint training but separate sharply in class-incremental (C-IL) settings (Van de Ven et al., 2022) (see Table 1). This motivates probing how negative-branch behavior affects plasticity under shift. Unless noted, all case studies use the same 4-layer CNN backbone, the Adam optimizer (Kingma, 2014), and training budget (full details in App. B), isolating activation effects from architectural or optimization confusion.
@@
i.i.d.
72.11
72.00
73.71
71.43
73.16
72.51
72.66
72.64
72.12
66.49
58.78
C-IL
24.41
28.57
32.95
22.71
24.43
20.91
22.79
27.59
27.49
26.44
25.47
Table 1: Values for Split-CIFAR-100 are reported as average accuracy (standard deviation) for 5 independent runs with identical architecture, optimizer, and budget. Performance differences across activation functions are modest under i.i.d. joint training but widen under class-incremental learning (C-IL). RReLU attains the top mean in both settings; the C-IL improvement is statistically significant (all $p<0.05$), whereas i.i.d. differences are not.
@@ 3 Case Study 1: Negative-Slope ‘Goldilocks Zone’
‘Goldilocks zone’ for negative slopes (with shape-matched comparison). Performance of activations with constant negative-branch slope (Leaky-ReLU, RReLU), reliably peaks for a moderate leak $0.6\!\lesssim\!\bar{s}\!\lesssim\!0.9$ and degrades once $\bar{s}\!\gtrsim\!0.9$ (Fig. 1A, Tab. 1), suggesting a ‘Goldilocks zone’ for negative slopes. To compare smooth-tailed functions on a common scale, we project their negative-branch behavior onto an effective slope axis, $\bar{s}=\mathbb{E}_{x<0}[\varphi^{\prime}(x)]$, representing the average derivative for negative inputs. When matched by $\bar{s}$, these smooth tails still underperform linear leaks within the ‘Goldilocks zone’ and exhibit higher dead-unit fractions (Fig. 1B), approaching (but not surpassing) the linear-leak peak only for $\bar{s}>1$.
Failure Modes for Slope Magnitude. We identify two distinct mechanisms defining the zone’s boundaries. As $\bar{s}\!\to\!0$, a dead-unit regime dominates ($\approx\!45\%$ inactive). We implement dead-unit here as effective inactivity, identifying units either stuck in saturation plateaus or contributing negligible magnitude relative to the layer’s scale. This inactivity strongly correlates with accuracy loss (Pearson r=0.51r{=}-0.51, $p{=}8.2{\times}10^{-28}$), and weakly but also significant with respect to final scaled gradient norms (Pearson r=0.11r{=}-0.11, p=0.029p{=}0.029) (App. C.1). Conversely, as $\bar{s}\!\to\!1$, performance degrades despite minimal dead units. Fig. 1C-D reveals this coincides with optimization instability: sharp spikes in principal curvature ($\lambda_{\max}$) and effective rank (App. B.7). Thus, sustaining plasticity requires a trade-off: avoiding gradient starvation (low $\bar{s}$) without inducing landscape stiffness (high $\bar{s}$).
Figure 1: A: Final accuracy vs. effective negative slope $\bar{s}$. B: Dead-unit fraction vs. $\bar{s}$. Linear-leak families peak for $\bar{s}\!\in\![0.6,0.9]$. Smooth-tailed activations are plotted on the same $\bar{s}$ axis; they underperform within the ‘Goldilocks zone’ and only approach the linear-leak peak when $\bar{s}\!>\!1$, reflecting concentrated near-zero responsiveness and vanishing tails. C: Effective rank of the gradient Gram matrix. D: Dominant $\lambda_{\max}$. Smooth-tailed activations show spikes at large $\bar{s}$, while constant-slope leaks remain comparatively stable.
Current adaptive, learnable slopes fail and need constraints to stay in-band. Under non-stationarity, the “right” negative-side responsiveness is not uniform across units or tasks. We therefore asked whether learnable or randomized slopes can find and maintain their specific ‘Goldilocks zone’. Per-neuron PReLU (PReLU-N) drifts below the band ($\approx 0.3-0.6$) over training and attains 30.1% ACC (see Figs.. C.2); layer/global scopes drift even further (Figs. C3). Thus, adaptivity is relevant—it offers robustness when a single fixed leak cannot serve all units—but unconstrained adaptation does not reliably remain in the pre-defined ‘Goldilocks zone’, which explains why learned slopes may sit outside it despite good (but suboptimal) ACC.
@@ 4 Case Study 2: Desaturation Dynamics Under Shocks
Every $C_{\ell}{=}10$ epochs we apply a one-epoch scaling shock by multiplying all pre-activations $z$ by $\gamma\in\{1.5,\,0.5,\,0.25,\,2.0\}$, then revert to $\gamma{=}1$ (App. D.1). Large $\gamma$ (e.g., $2.0$) pushes $z$ into negative tails or positive plateaus (saturation); small $\gamma$ (e.g., $0.5$) produces the mirrored event. Each activation uses its best negative-slope setting from Case Study 3 (Table B2). A unit is saturated at a step if $|\varphi^{\prime}(x)|<10^{-3}$. We report: (i) Peak SF: the maximum saturated fraction immediately after each shock; (ii) AUSC: the area under the saturation curve over the recovery window (lower is better); (iii) Recovery time $\bm{\tau_{95}}$: steps needed to regain $95\%$ of pre-shock performance (App. D.4).
@@
The width of the dead band predicts shock sensitivity. Beyond the findings of derivative floor and sidedness, we ask how much of the input range of an activation produces nearly zero gradients. We define a Dead-Band Width ($DBW$) as the fraction of a typical pre-activation range (e.g., [-100, 100]) where the magnitude of the activation’s first derivative, $\lvert\varphi^{\prime}(x)\rvert$ falls below certain threshold, $\epsilon<10^{-3}$. Analytically computed $DBW$, $|\varphi^{\prime}(x)|<10^{-3}$, strongly tracks desaturation outcomes across activations.
We hypothesize that this analytically derived $DBW$ will positively correlate with experimentally observed adverse saturation dynamics. Intuitively, a wider intrinsic dead-band means pre-activation scaling shocks are more likely to push numerous units into these unresponsive, vanishing-gradient regions and keep them there. This, in turn, diminishes the network’s ability to desaturate, which we expect to manifest as a greater overall saturation impact (higher AUSC) and increased non-recovery rates.
Figure 4 strongly supports this hypothesis, revealing that the analytical Dead-Band Width Score is strongly and significantly correlated with adverse saturation outcomes. Activations with a higher $DBW$ experience a greater overall saturation impact (Avg. AUSC, r=0.81,p=0.0016r=0.81,\;p=0.0016) and are much more likely to fail recovery entirely (Avg. SF Non-Recovery Rate, r=0.84,p=0.0013r=0.84,\;p=0.0013). DBWDBW does not predict recovery speed. Successful desaturation, when it occurs, is consistently fast (around one epoch) and shows no correlation with the $DBW$. This suggests the $DBW$ score predicts the likelihood and severity of saturation, but not the speed of recovery.
Figure 4: Correlation of Dead-Band Width Score with Saturation Recovery Metrics (All Gammas Aggregated). (Left): Average Area Under Saturation Curve (Avg. AUSC) vs. Dead-Band Width Score. A strong positive correlation (Pearson r=0.81,p=0.0016r=0.81,\;p=0.0016) is observed. (Middle): Average Saturation Fraction (SF) Recovery Time (for successful recoveries, measured by epochs) vs. Dead-Band Width Score. No significant correlation is found (Pearson r=0.25,p=0.45r=-0.25,\;p=0.45). (Right): Average SF Non-Recovery Rate (%) vs. Dead-Band Width Score. A strong positive correlation (Pearson r=0.84,p=0.0013r=0.84,\;p=0.0013) is observed, indicating functions more prone to saturation are more likely to fail SF recovery.
@@ 5.1 Smooth-Leaky and Randomized Smooth-Leaky
Figure 5: Smooth-Leaky with $\alpha{=}0.1$, p=3.0p{=}3.0, c=5.0c{=}5.0. Randomized Smooth-Leaky draws $\alpha$ from bounds.
@@
This randomized variant preserves the strict floor and $C^{1}$ transition while encouraging robustness to small variations in negative-side responsiveness. Limitations of multi-parameter design are explained in App. B.2.
@@ 6 Continual Supervised Learning
ReLU 78.85  $\pm$ 0.06 20.03  $\pm$ 2.46 25.79  $\pm$ 6.18 4.76  $\pm$ 1.01 73.71  $\pm$ 0.43
Leaky-ReLU 84.14  $\pm$ 0.01 91.53  $\pm$ 0.18 98.34  $\pm$ 0.01 48.86  $\pm$ 0.70 85.28  $\pm$ 0.20
Sigmoid 76.96  $\pm$ 0.07 79.59  $\pm$ 0.75 52.24  $\pm$ 2.99 1.79  $\pm$ 0.19 63.89  $\pm$ 7.38
Tanh 70.32  $\pm$ 0.54 63.40  $\pm$ 0.12 58.56  $\pm$ 1.05 28.59  $\pm$ 2.34 70.97  $\pm$ 0.44
RReLU 83.95  $\pm$ 0.02 93.10  $\pm$ 0.02 98.02  $\pm$ 0.03 53.60  $\pm$ 1.06 84.97  $\pm$ 0.17
PReLU 82.62  $\pm$ 0.05 92.67  $\pm$ 0.23 96.86  $\pm$ 0.32 43.30  $\pm$ 0.61 82.37  $\pm$ 0.11
Swish (SiLU) 83.41  $\pm$ 0.03 67.73  $\pm$ 0.46 87.40  $\pm$ 2.42 35.31  $\pm$ 1.87 82.64  $\pm$ 0.99
GeLU 78.97  $\pm$ 0.09 38.79  $\pm$ 0.95 42.85  $\pm$ 2.12 17.60  $\pm$ 1.71 75.49  $\pm$ 0.11
CeLU 82.93  $\pm$ 0.04 37.16  $\pm$ 0.90 29.64  $\pm$ 10.44 54.23  $\pm$ 1.44 81.15  $\pm$ 0.68
eLU 80.50  $\pm$ 0.09 84.23  $\pm$ 0.70 57.45  $\pm$ 20.16 47.64  $\pm$ 1.44 80.10  $\pm$ 0.34
SeLU 80.43  $\pm$ 0.16 79.95  $\pm$ 0.91 84.61  $\pm$ 2.07 49.07  $\pm$ 1.25 80.98  $\pm$ 0.49
CReLU 82.66  $\pm$ 0.04 89.47  $\pm$ 0.28 92.90  $\pm$ 0.13 20.56  $\pm$ 2.28 84.85  $\pm$ 0.25
Rational 80.08  $\pm$ 0.05 92.35  $\pm$ 1.97 94.82  $\pm$ 0.75 40.41  $\pm$ 4.21 80.65  $\pm$ 0.38
SwiGLU 77.69  $\pm$ 0.26 31.20  $\pm$ 2.10 83.06  $\pm$ 3.51 9.57  $\pm$ 1.81 63.57  $\pm$ 2.04
Deep Fourier 83.69  $\pm$ 0.04 92.61  $\pm$ 0.04 96.24  $\pm$ 0.51 72.29  $\pm$ 2.11 76.03  $\pm$ 0.75
Smooth-Leaky 84.03  $\pm$ 0.02 91.69  $\pm$ 0.12 98.36  $\pm$ 0.00 49.87  $\pm$ 1.67 85.38  $\pm$ 0.25
Rand. Smooth-Leaky 84.26  $\pm$ 0.02 93.33  $\pm$ 0.05 98.42  $\pm$ 0.01 57.01  $\pm$ 1.59 86.23  $\pm$ 0.13
Table 2: Total Average Online Task Accuracy (%) on Continual Supervised Benchmarks, averaged over 5 independent runs. Values are reported as mean $\pm$ standard deviation (SD). Statistical significance between the top two performers in each column was determined using an independent two-sample Welch’s t-test (p << 0.05). Rand. Smooth-Leaky is statistically significant with respect to the next best-performing activation (Smooth-Leaky). Smooth-Leaky is also significant compared to the next best in Rand. Label CIFAR, CIFAR 5+1 and Continual ImageNet.
Across the five continual benchmarks, we observe a clear pattern, reported in Tab. 2. First, rectifiers with a learnable or randomized negative branch dominate: Leaky-ReLU, RReLU, PReLU, Smooth-Leaky, and Rand. Smooth-Leaky consistently outperform ReLU—especially on the harder settings—while smooth rectifiers such as Swish/SiLU are competitive but typically trail the best leaky-family members. Second, we again observe the first reported ‘Goldilocks zone’ for the negative branch (cf. Sec. 3): the strongest performers cluster around an initial/effective negative slope in the range [0.6, 0.9][0.6,\,0.9] (including the mean of RReLU bounds and neuron-wise $\alpha$ in PReLU). We also evaluated CReLU and Rational activations  Abbas et al. (2023); Delfosse et al. (2021b), Deep Fourier Features Lewandowski et al. (2024), which satisfies our criteria for a smooth, non-zero gradient floor, maintaining responsiveness via periodic oscillation rather than a fixed linear slope, and SwiGLU Shazeer (2020), a gated nonlinearity that that gained recent popularity due to improve Transformer feed-forward layers and widely adopted in PaLM and LLaMA Touvron et al. (2023); Chowdhery et al. (2023). Constrained rationals were excluded, as previous work shows that they improve RL stability but reduce plasticity Surdej et al. (2025), which is our main focus. In our experiments, CReLU, Rational and Deep Fourier outperform ReLU, but remain below other standard activations and our proposed variants (see Tab. E2 and Fig. E1).